# Quantum simulation of many-body dynamics with noise-robust Trotter decomposition based on symmetric structures

### v2

@@ III Proposed Decomposition
Using the circuit implementation of $\exp(-{\mathrm{i}}\hat{O}^{\prime}_{0}\Delta t^{\prime})$, the whole quantum circuit of the proposed Trotter block for $\Delta t^{\prime}$ in Eq. (10) becomes Fig. 5(a), which contains $14$ CNOT gates in each four-qubit Trotter block. Thus, our circuit construction consumes on average $3.5$ CNOT gates per qubit in a single Trotter block with evolution time $\Delta t^{\prime}$. Moreover, taking the ratio between $\Delta t$ and $\Delta t^{\prime}$ into account, our circuit construction requires only $3.5m=2.625n$ CNOT gates to perform $t=n\Delta t$ time evolution for. Since the conventional Trotter circuit requires an average of $3n$ CNOT gates per qubit up to time $t$, yielding a reduction rate of $2.625n/3n=0.875$ compared to the original Trotter blocks. This reduction of CNOT gates significantly contributes to the noise resilience of our proposed Trotter decomposition.
Figure 6: Comparison between the conventional Trotter decomposition (Fig. 1) and the proposed decomposition (Fig. 5), where plots of the conventional Trotter decomposition are colored red and those of the proposed decomposition are colored blue. (a) The state infidelity of the evolved state under the depolarizing noise with $p_{1}=1.0\times 10^{-4}$ to the ideally evolved state without noise, scaling with different numbers of Trotter steps $n$. Note that the x-axis for the proposed method is normalized to $3m/4$ on the plot so that the residual error of Eq. (3) and Eq. (10) become compatible. (b) The state infidelity of the evolved state under the noise to the ideally evolved state without noise, scaling with different depolarizing probabilities $p_{1}\in\{1.0\times 10^{-5},3.0\times 10^{-4},1.0\times 10^{-4},3.0\times 10^{-4}\}$ and $p_{2}=10p_{1}$.
@@ IV Experiments
To demonstrate the practicality of the proposed new Trotter decomposition, we simulate the time evolution of the $XXX$ Heisenberg model and calculate the fidelity $F(\tilde{\rho},\rho_{\mathrm{ideal}})=\operatorname{Tr}\left[\left(\rho_{\mathrm{ideal}}^{1/2}\tilde{\rho}\rho_{\mathrm{ideal}}^{1/2}\right)^{1/2}\right]^{2}$ of the resulting state $\tilde{\rho}$ to the ideally evolved state $\rho_{\mathrm{ideal}}$. We use the noisy simulator, real-device emulator fake_jakarta, and real quantum device ibmq_jakarta provided by IBM Quantum Platform.
IV.1 Numerical simulation under depolarizing noise
Figure 7: (a) The number of the CNOT gates scaling with different numbers of Trotter steps $n$. Note that the x-axis for the proposed method is normalized to $3m/4$ on the plot so that the residual error of Eq. (3) and Eq. (10) become compatible. (b) The infidelity between the evolved state under the noise and the ideally evolved state without noise, scaling with different Trotter steps.
Figure 8: (a) The infidelity of the evolved state to the expected noise-free state, simulated by the two encoding-decoding strategies among different Trotter steps $m$. (b) The infidelity of the evolved state to the expected noise-free state, for different QEM levels among different Trotter steps $m$.
@@ IV.1 Numerical simulation under depolarizing noise
We first perform the noisy numerical simulation of the nine-site $XXX$ Heisenberg model up to a fixed evolution time $t=\pi$. We introduce depolarizing noise in both single-qubit gates and two-qubit gates with depolarizing probabilities $p_{1}=1.0\times 10^{-4}$ and $p_{2}=10p_{1}=1.0\times 10^{-3}$, respectively, which reflect the typical noise levels of the current quantum hardware. Starting from an arbitrary initial state, which we set to $|110100110\rangle$, we compare the state fidelity between the conventional Trotter circuit in Fig. 1 and the proposed Trotter circuit in Fig. 5 with different Trotter steps among $n,m\in\{4,8,12,\ldots,96\}$.
@@
The simulated results between the conventional Trotter decomposition and the proposed decomposition are plotted in Fig. 6. We observe from Fig. 6(a) that the proposed method achieves lower state infidelity than the conventional approach. We also observe that there exists an optimal number of Trotter steps that balances the simulation accuracy and the noise effect induced when increasing the Trotter steps. Focusing on the infidelity with the optimal number of Trotter steps, Fig. 6(b) visualizes the advantage of our proposed method over the conventional method in terms of lower infidelity among different noise levels.
The results mentioned above imply that our proposed decomposition is more noise-robust thanks to its reduced use of CNOT gates. Fig. 7(a) visualizes the number of CNOT gates used in the whole quantum circuit between the proposed method and the conventional one, reflecting the estimated reduction ratio $0.875$. We further plot the relation between the number of CNOT gates and the state infidelity in Fig. 7(b), where the proposed method achieves smaller state infidelities with the same number of CNOT gates. This implies that the proposed quantum circuit constructs more noise-resilient Trotter blocks than the conventional circuit.
IV.2 Real-device experiments with error suppression
@@
Based on the above, we compute the infidelity of the resulting state at the evolution time $t=\pi$ evolved from a given initial state $|110\rangle$ at $t=0$. We add quantum error mitigation (QEM) Li and Benjamin (2017); Temme, Bravyi, and Gambetta (2017); Endo, Benjamin, and Li (2018); Giurgica-Tiron et al. (2020); Yang, Raymond, and Uno (2022); Yang et al. (2023); Cai et al. (2023); Endo et al. (2021) to reduce the noise effect through classical post-processing. Particularly, we use quantum readout error mitigation (QREM) Yang, Raymond, and Uno (2022) and zero-noise extrapolation (ZNE) Temme, Bravyi, and Gambetta (2017); Giurgica-Tiron et al. (2020). We use the digital ZNE method Giurgica-Tiron et al. (2020) with the linear fitting method and the scale factors $1.0$, $2.0$ and $3.0$, provided by Mitiq LaRose et al. (2022). Furthermore, the Pauli twirling technique Bennett et al. (1996); Knill (2004); Wallman and Emerson (2016) is also combined with ZNE, referring to the implementation by Berthusen et al. Berthusen et al. (2021). Each quantum circuit is executed with $8192$ shots, and the infidelity is averaged over $8$ samples.
First, we examine the performance among Trotter steps $\{4,5,6,7,8,9,10,20,30,40\}$ on the noisy simulator fake_jakarta, comparing the two encoding-decoding methodologies: the general encoder $\hat{U}_{\mathrm{enc}}$ and general decoder $\hat{U}_{\mathrm{enc}}^{\dagger}$ (general-general), and the aforementioned shallow encoding and specific encoding (shallow-specific). Here, we apply only QREM to noisy results before computing the infidelity. The result is shown in Fig. 8(a), where both encoding-decoding methods achieve the infidelity below $0.2$ with more than $10$ Trotter steps and the shallow-specific method further achieves the infidelity smaller than $0.1$.
The effect of QEM methods is also investigated under fake_jakarta. Here, we set the configuration to shallow encoding and specific decoding. We observe from Fig. 8(b) that both QREM and ZNE contribute to reducing the infidelity. The instability of ZNE can be improved by adding Pauli twirling that tailors the noise to a stochastic Pauli channel. We also see that combining Pauli twirling further enhances the accuracy gains achieved by finer Trotter step decomposition.
@@
Finally, we examine the time evolution from $t=0$ to $t=\pi$ on the real quantum device ibmq_jakarta. We execute $100$ Trotter steps with and without QEM under the two encoding-decoding methods. The state fidelity is then calculated by reconstructing the density matrix through state tomography Raymer, Beck, and McAlister (1994); Leonhardt (1995); Leibfried et al. (1996).
@@ V Conclusion
This work opens up several promising directions. First, thanks to the generality of the proposed approach, it can be extended to a broader range of physical and chemical models, such as transverse-field Ising models and Fermi-Hubbard models. Next, integrating more recent QEM techniques based on ZNE Endo et al. (2019); Hakkaku et al. (2025) would further suppress both coherent and algorithmic errors in Trotterized circuits. Moreover, resource-efficient implementations using subspace expansion methods Yoshioka et al. (2022); Yang et al. (2023) can also be adapted to our framework to mitigate both coherent and stochastic errors. In combination with quantum–classical divide-and-conquer approaches Sun et al. (2022); Eddins et al. (2022); Yuan et al. (2021); Harada et al. (2025), these techniques would further enhance the practical feasibility of our method on near-term quantum devices.
@@ Acknowledgements.
The result of real-device experiments is obtained as a solution to the IBM Quantum Awards: Open Science Prize 2021, for which the initial manuscript has been publicly available at Yang and Negishi (2022) since April 2022. B.Y. and N.N. sincerely thank all those who made this contest possible.
@@
The code and data used in the experiments in this work are publicly available on the GitHub page: https://github.com/BOBO1997/osp_solutions Yang and Negishi (2022).

### v3

@@ III Proposed Decomposition
Using the circuit implementation of $\exp\left(-{\mathrm{i}}\hat{A}_{k}\Delta t^{\prime}\right)$ and $\exp\left(-{\mathrm{i}}\hat{B}_{k}\Delta t^{\prime}\right)$, the whole quantum circuit of the proposed Trotter block for $\Delta t^{\prime}$ in Eq. (13) is then described by Fig. 7. The single Trotter block requires $8\times 4=32$ CNOT gates per four qubits. Moreover, by considering the offsets between the two CNOT gates in the red-dashed rectangles in Fig.7, we can finally reduce CNOT gates to $28$ in each four-qubit Trotter block. Thus, our circuit construction consumes on average seven CNOT gates per qubit in a single Trotter block with evolution time $2\Delta t^{\prime}$. Moreover, taking the ratio between $\Delta t$ and $\Delta t^{\prime}$ into account, our circuit construction requires only $7m=1.75n$ CNOT gates to perform $t=n\Delta t$ time evolution for. Since the conventional Trotter circuit requires an average of $3n$ CNOT gates per qubit up to time $t$, yielding a reduction rate of $1.75n/3n=0.583$ compared to the original Trotter blocks. This reduction of CNOT gates significantly contributes to the noise resilience of our proposed Trotter decomposition.
@@ IV Experiments
To demonstrate the practicality of the proposed Trotter decomposition, we simulate the time evolution of the $XXX$ Heisenberg model and calculate the fidelity $F(\tilde{\rho},\rho_{\mathrm{ideal}})=\operatorname{Tr}\left[\left(\rho_{\mathrm{ideal}}^{1/2}\tilde{\rho}\rho_{\mathrm{ideal}}^{1/2}\right)^{1/2}\right]^{2}$ of the resulting state $\tilde{\rho}$ to the ideally evolved state $\rho_{\mathrm{ideal}}$.
We also compare the process fidelity to the accurate time evolution from $t=0$ to $t=\pi$ between the conventional Trotter unitary operation and the proposed Trotter unitary operation. Note that the process fidelitynielsen00 is defined as the state fidelity between Choi matrices of the two completely positive trace-preserving (CPTP) maps of interest. In the following experiments, we use the noisy and noise-free density matrix simulator, real-device emulator fake_jakarta, and real quantum device ibmq_jakarta provided by IBM Quantum Platform.
IV.1 Noise-free process fidelity
We first examine the advantage in the convergence rate of approximation error of the proposed Trotter decompositions over the conventional one without the initial state dependency. To see this, we compare the process fidelity of the conventional and of the proposed approaches to the theoretically predicted time evolution without noise. Here, we use nine-site $XXX$ Heisenberg model up to a fixed evolution time $t=\pi$. For the conventional Trotter circuit (Fig. 2) we perform $19$ different Trotter iterations among $n\in\{40,60,80,\ldots,400\}$, and for the proposed Trotter circuit (Fig. 7), among $m\in\{20,30,40,\ldots,200\}$.
The results are shown in Fig. 8, where we plot the process infidelity of the conventional method and the proposed method. Figure 8(a) compares the process infidelity with different Trotter iterations where both methods use the same time intervals $\Delta t=\Delta t^{\prime}$. Figure 8(b) compares the process infidelity on the same basis of the number of the total CNOT gates in the Trotter circuit. We observe that the process infidelity of the proposed method seems to be halved from that of the conventional one in Fig. 8(a), which is consistent with theoretical prediction under the setup of $\Delta t=\Delta t^{\prime}$ and $m=n/2$. Besides, Fig. 8(b) implies that the proposed method is more noise-robust in the sense that it uses fewer CNOT gates which are considered to be the main source of noise on the current quantum hardware.
IV.2 State fidelity under depolarizing noise
Figure 9: The state infidelity of noisy time evolution starting from the Néel state $|101010101\rangle$ to the ideally evolved state. Plots of the conventional Trotter decomposition (Fig. 2) are colored red and those of the proposed decomposition (Fig. 7) are colored blue. (a) The state infidelity of the evolved state under the depolarizing noise with $p_{1}=1.0\times 10^{-6}$ to the ideally evolved state, scaling with the number of Trotter iterations. The x-axis $k$ denotes the $k$-th iteration of $n\in\{4,12,20,\ldots,196\}$ for the conventional method and $m\in\{2,6,10,\ldots,98\}$ for the proposed method to make the time intervals $\Delta t$ and $\Delta t^{\prime}$ consistent between each method, i.e. $\Delta t=\Delta t^{\prime}$. (b) The state infidelity of the evolved state under the depolarizing noise with $p_{1}=1.0\times 10^{-6}$ to the ideally evolved state, scaling with the number of CNOT gates.
Figure 10: The minimum state infidelity of the evolved state under the noise with the conventional and proposed approaches to the ideally evolved state without noise, scaling with different depolarizing probabilities $p_{1}\in\{1.0\times 10^{-7},3.0\times 10^{-7},1.0\times 10^{-6},3.0\times 10^{-6},1.0\times 10^{-5}\}$ and $p_{2}=10p_{1}$.
We next perform the noisy numerical simulation of the nine-site $XXX$ Heisenberg model up to a fixed evolution time $t=\pi$.
@@ IV.2 State fidelity under depolarizing noise
We introduce depolarizing noise in both single-qubit gates and two-qubit gates with depolarizing probabilities $p_{1}=1.0\times 10^{-6}$ and $p_{2}=10p_{1}=1.0\times 10^{-5}$, respectively, which reflect the noise levels of the current and near-future quantum hardware Hughes2025Trapped-ion; Ransford2025Helios. Starting from the Néel state $|101010101\rangle$, we compare the state fidelity between the conventional approach for $25$ different Trotter iterations among $n\in\{4,12,20,\ldots,196\}$ and the proposed approach for $25$ different Trotter iterations among $m\in\{2,6,10,\ldots,98\}$.
@@
The simulated results between the conventional Trotter decomposition and the proposed decomposition are plotted in Fig. 9. We observe that there exists an optimal number of Trotter iterations that balances the simulation accuracy and the noise effect induced when increasing the Trotter iterations. From Fig. 9(a), we see that the optimal state infidelity by the proposed method achieves lower state infidelity than the conventional approach.
Note that the cross of the plots between the conventional method and the proposed method when $k=12$ in Fig. 9(a) results from the condition to make $\Delta t=\Delta t^{\prime}$ in the x-axis, where the extent of the residual error by the proposed method is designed to be the half of the residual error by the conventional method. This means that the proposed method uses correspondingly deeper circuits with more CNOT gates to achieve such a better residual error. Under the noisy execution, this appears as the cross of the plots. When adjusting the x-axis relative to the number of CNOT gates in Fig. 9(b), we then see that the performance of the proposed method is better than the conventional one for larger Trotter iterations. This implies that our proposed decomposition is more noise-robust to achieve a higher state infidelity under the noisy execution. We here also remark that the conventional method outperforms the proposed method for smaller Trotter iterations, which can be explained by insufficient precision to approximate $\exp\left(-\mathrm{i}\hat{H}_{\mathrm{eff}}\Delta t^{\prime}\right)$ in Eq. (21).
Focusing on the infidelity with the optimal number of Trotter iterations when simulating time evolution starting from the Néel state, Fig. 10 visualizes the advantage of our proposed method over the conventional method in terms of lower infidelity under different noise levels ranging among $p_{1}\in\{1.0\times 10^{-7},3.0\times 10^{-7},1.0\times 10^{-6},3.0\times 10^{-6},1.0\times 10^{-5}\}$ and $p_{2}=10p_{1}$. Note that the latest quantum devices have already reach the regime where the single-qubit and two-qubit gate error probability take $10^{-5}$ and $10^{-4}$, respectively Hughes2025Trapped-ion; Ransford2025Helios. Observing from Fig. 10, the advantage of the proposed method becomes clearer when the error rates become smaller. This means that our approach would further outperform the conventional approach in the near-future quantum devices with smaller noise levels.
Figure 11: (a) The state infidelity of the evolved state to the expected noise-free state, simulated by the two encoding-decoding strategies among different Trotter iterations $m$. (b) The state infidelity of the evolved state to the expected noise-free state, for different QEM levels among different Trotter iterations $m$.
IV.3 Real-device experiments with error suppression
@@
In implementing the proposed Trotter decomposition on real quantum hardware ibmq_jakarta and its noise-calibrated simulator fake_jakarta, we simulate the time evolution of the three-site $XXX$ Heisenberg model from $t=0$ to $t=\pi$. This real device experiments with $N=3$ corresponds to using only the propagator $\exp\left(-{\mathrm{i}}\hat{A}_{1}\Delta t^{\prime}\right)$ in Eq. (14), which can be seen as the simplest case of the proposed method.
@@
Based on the above, we compute the infidelity of the resulting state at the evolution time $t=\pi$ evolved from a given initial state $|110\rangle$ at $t=0$. We add quantum error mitigation (QEM) li2017efficient; Temme_2017; endo2018practical; digital_zne; yang2022efficient; Yang2025Resource-Efficient; cai2023quantum; endo2021hybrid to reduce the noise effect through classical post-processing. Particularly, we use quantum readout error mitigation (QREM) yang2022efficient and zero-noise extrapolation (ZNE) Temme_2017; digital_zne. We use the digital ZNE method digital_zne with the linear fitting method and the scale factors $1.0$, $2.0$ and $3.0$, provided by Mitiq mitiq. Furthermore, the Pauli twirling technique Bennett1996Purification; Knill2004Fault-Tolerant; Wallman2016Noise is also combined with ZNE, referring to the implementation by Berthusen et al. Berthusen_2021. Each quantum circuit is executed with $8192$ shots, and the infidelity is averaged over $8$ samples.
First, we examine the performance among Trotter iterations $\{4,5,6,7,8,9,10,20,30,40\}$ on the noisy simulator fake_jakarta, comparing the two encoding-decoding methodologies: the general encoder $\hat{U}_{\mathrm{enc}}$ and general decoder $\hat{U}_{\mathrm{enc}}^{\dagger}$ (general-general), and the aforementioned shallow encoding and specific encoding (shallow-specific). Here, we apply only QREM to noisy results before computing the infidelity. The result is shown in Fig. 11(a), where both encoding-decoding methods achieve the infidelity below $0.2$ with more than $10$ Trotter iterations and the shallow-specific method further achieves the infidelity smaller than $0.1$.
The effect of QEM methods is also investigated under fake_jakarta. Here, we set the configuration to shallow encoding and specific decoding. We observe from Fig. 11(b) that both QREM and ZNE contribute to reducing the infidelity. The instability of ZNE can be improved by adding Pauli twirling that tailors the noise to a stochastic Pauli channel. We also see that combining Pauli twirling further enhances the accuracy gains achieved by finer Trotter decomposition.
@@
Finally, we examine the time evolution from $t=0$ to $t=\pi$ on the real quantum device ibmq_jakarta. We execute $100$ Trotter iterations with and without QEM under the two encoding-decoding methods. The state fidelity is then calculated by reconstructing the density matrix through state tomography Raymer1994Complex; Leonhardt1995Quantum-State; Leibfried1996Experimental.
@@ V Conclusion
This work opens up several promising directions. First, owing to the generality of the proposed approach, it can be extended to a broader class of physical and chemical models. In particular, our framework applies to models that share the same underlying symmetric structure as the Heisenberg model after an equivalent mapping, such as the Jordan-Wigner transformation JWT1; JWT2, which maps fermionic creation and annihilation operators to Pauli operators. Through such transformations, electronic quantum states would be able be simulated within our present framework. A systematic characterization of the classes of physical systems that can be equivalently mapped onto this symmetric form is left as a significant direction for future work.
Next, integrating more recent QEM techniques based on ZNE Endo2019Mitigating; Hakkaku2025Data-Efficient would further suppress both coherent and algorithmic errors in Trotterized circuits. Moreover, resource-efficient implementations using subspace expansion methods Yoshioka2022Generalized; Yang2025Resource-Efficient can also be adapted to our framework to mitigate both coherent and stochastic errors. In combination with quantum–classical divide-and-conquer approaches Sun2022Perturbative; Eddins2022Doubling; Yuan2021Quantum; Harada2025densitymatrix, these techniques would further enhance the practical feasibility of our method on near-term quantum devices.
@@ Acknowledgements.
The result of real-device experiments is obtained as a solution to the IBM Quantum Awards: Open Science Prize 2021, for which the initial manuscript has been publicly available at github_osp_solutions since April 2022. B.Y. and N.N. sincerely thank all those who made this contest possible.
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Bo Yang: Conceptualization (equal); Data curation (lead); Formal analysis (supporting); Funding acquisition (equal); Investigation (lead); Methodology (supporting); Project administration (equal); Resources (equal); Software (lead); Supervision (equal); Validation (equal); Visualization (equal); Writing - original draft (equal); Writing - review & editing (equal).
Naoki Negishi: Conceptualization (equal); Data curation (supporting); Formal analysis (lead); Funding acquisition (equal); Investigation (supporting); Methodology (lead); Project administration (equal); Resources (equal); Software (supporting); Supervision (equal); Validation (equal); Visualization (equal); Writing - original draft (equal); Writing - review & editing (equal).
@@
The code and data used in the experiments in this work are publicly available on the GitHub page: https://github.com/BOBO1997/osp_solutions github_osp_solutions.